how many pounds of coffee worth 10$ a pound should be added to 20 pounds of coffee worth 4$ a pound to get a mixture worth 5$ a pound
Answer by ewatrrr(24785)

x how many pounds of coffee worth 10$ a pound that should be added
10x + 4(20lb) = 5(20lb + x)
x = (20lb)/5 = 4lb, pounds of coffee worth 10$ a pound that should be added